# Yangqin Jiang, Chao Huang

### v1

@@ Abstract
With the advancement of multimodal large language models (MLLMs), building GUI agent systems has become an increasingly promising direction—especially for mobile platforms, given their rich app ecosystems and intuitive touch interactions. Yet mobile GUI agents face a critical dilemma: truly on-device models (4B or smaller) lack sufficient performance, while capable models (starting from 7B) are either too large for mobile deployment or prohibitively costly (e.g., cloud-only closed-source MLLMs). To resolve this, we propose LightAgent, a mobile agentic foundation model solution that leverages device-cloud collaboration to tap the cost-efficiency of on-device models and the high capability of cloud models, while avoiding their drawbacks. Specifically, LightAgent enhances Qwen2.5-VL-3B via two-stage SFT→GRPO training on synthetic GUI data for strong decision-making, integrates an efficient long-reasoning mechanism to utilize historical interactions under tight resources, and defaults to on-device execution—only escalating challenging subtasks to the cloud via real-time complexity assessment. Experiments on the online AndroidLab benchmark and diverse apps show LightAgent matches or nears larger models, with a significant reduction in cloud costs. We have made our LightAgent available at: https://github.com/HKUDS/LightAgent.
@@ 1 Introduction
The growing capability of multimodal large language models (MLLMs) enables AI agents to perceive and act within visual environments, particularly through Graphical User Interfaces (GUIs) (Wu et al., 2024b; Qi et al., 2024). Mobile platforms offer a promising domain for this technology for two reasons: first, their vast app ecosystems provide a realistic and diverse testbed, and second, their touchscreen interactions are limited to an intuitive set of primitives, resulting in a compact action space. Despite these advantages, mobile platforms introduce distinct challenges, chiefly severe computational and memory limitations. In light of these factors, our goal is to develop an effective mobile GUI agent, viewing it as a practical milestone toward general-purpose AI.
Current research falls broadly into two groups. The first involves targeted training of open-source MLLMs specifically for GUI-related tasks (Qin et al., 2025; Dai et al., 2025; Liu et al., 2024). These models are relatively compact in size and have achieved notable progress; for instance, UI-Tars-7B (Qin et al., 2025) outperforms the larger Qwen2.5-VL-32B (Bai et al., 2025) on mobile GUI tasks. The second approach leverages general-purpose closed-source MLLMs by constructing multi-agent systems and designing well-structured execution pipelines. Thanks to the powerful multimodal comprehension capabilities of state-of-the-art (SOTA) closed-source MLLMs—such as GPT-5 (OpenAI, 2025), Claude-Sonnet-4 (Anthropic, 2025), and Gemini-2.5-Pro (Comanici et al., 2025)—their performance on mobile GUI tasks can even surpass that of models specifically trained for such tasks. However, there is no free lunch. Although the performance of models like UI-Tars-7B is impressive given their 7B scale, MLLMs of this size still impose a prohibitive computational burden on contemporary smartphones (Laskaridis et al., 2024). A more practical 2B–3B scale, by contrast, typically yields MLLMs with limited capabilities (Lin et al., 2025). On the other hand, multi-agent systems based on advanced closed-source MLLMs are plagued by high costs. SOTA closed-source MLLMs are often expensive, and spending several to dozens of dollars to complete a single mobile task is financially impractical.
@@
To answer to the aforementioned questions, we introduce LightAgent: a mobile agentic foundation models solution combining a lightweight device–cloud collaborative GUI agent framework and a 3B-parameter on-device GUI agent model. Specifically, for Question 1, based on a lightweight open-source MLLM (i.e., Qwen2.5-VL-3B), we employ a synthetic data generation pipeline and conduct two-stage training comprising supervised fine-tuning (SFT) and group relative policy optimization (GRPO), resulting in a compact yet powerful on-device GUI agent that achieves performance comparable to larger-scale models. Concurrently, we design an efficient long-reasoning GUI agent paradigm, which is capable of effectively processing historical operation information and endowing the agent with reasoning capabilities to further enhance its performance.
@@
  • Device-cloud collaborative agent system. We propose a device-cloud collaborative agent system that dynamically orchestrates tasks between on-device and cloud models via real-time complexity assessment. This mechanism enables seamless switching to the cloud only when necessary, compensating for on-device limitations, achieving an optimal balance between high performance and significantly reduced operational costs.
  • Comprehensive Evaluation. We evaluate LightAgent through online experiments on the Android‑based AndroidLab benchmark, complemented by dozens of custom tasks across popular apps to assess real‑world performance. Furthermore, we conduct extensive ablation studies and investigate the impact of different fine‑tuning strategies on the lightweight MLLM’s capabilities.
2 Methodology
Figure 2: Overall framework of the proposed LightAgent.
To effectively address mobile GUI agent tasks, we propose the LightAgent framework, which has three key modules. Section 2.1 covers strategies to mitigate compact MLLMs’ challenges in GUI tasks—limited capacity and constrained context length. Section 2.2 details a device-cloud collaborative agent system that dynamically schedules on-device and cloud models by task difficulty, balancing cloud resource consumption with GUI task completion rate. Section 2.3 outlines ways to fully leverage limited training data to maximize small MLLMs’ performance gains on GUI tasks.
@@ 2.1 Efficient Reasoning GUI Agent
The on-device GUI agent faces two key challenges. First, mobile devices’ limited computing power necessitates small-sized MLLMs (e.g., Qwen2.5-VL-3B (Bai et al., 2025)), which currently lack sufficient performance for mobile GUI tasks. To tackle this, LightAgent enhances the GUI agent with extended chain-of-thought (CoT) reasoning (Wei et al., 2022) during testing, using test-time scaling laws (Snell et al., 2024) to boost its capability. Second, limited on-device resources impede long-context handling: high-resolution GUI images take up much available context length, and managing the agent’s execution history is challenging. To alleviate this, LightAgent uses an efficient text-based summarization scheme—compressing each step’s state into compact textual representations—to support the agent’s long historical context. The output format template is presented in Figure 3, and the detailed instruction template is provided in Appendix A.2.1.
@@
2.1.1 Long-Horizon Reasoning Enhancement
Mobile GUI tasks are often difficult and complex, and humans also use step-by-step reasoning for such operations. Motivated by the success of CoT reasoning and test-time scaling laws, it is natural to apply similar long-form reasoning enhancements to GUI agents.
Specifically, the GUI agent’s reasoning—encompassing all factors for task completion—follows a multi-step process: First, it analyzes actionable elements on the current interface and, using historical data, assesses if prior actions met their goals. Next, it evaluates progress toward the user’s task goal and identifies necessary follow-up actions. Lastly, it selects from available functions and their parameters to generate the final output. Notably, if historical data shows prior operations failed to yield expected results, the model proactively reflects and adjusts its approach—avoiding repeated errors and potential loops.
2.1.2 Efficient Memory Management
@@
To address this, LightAgent uses a textual summarization approach: at each step, it distills all information relevant to future actions. As shown in the <STATE_ASSESSMENT> field of Figure 3, these summaries include the current interface state, task progress, the agent’s inferred next action, expected post-action outcome, and potential issues. Much of this content comes from condensing the <REASONING> section—critical for effective stepwise reasoning, especially reflective error correction. Crucially, textual summaries use far fewer tokens than images, enabling long-term history retention (e.g., 10–20 steps) as contextual input, even in resource-constrained mobile environments.
@@ 2.1.3 Overall Process
with $\circ$ denoting sequence concatenation and $h_{0}=\epsilon$. The process repeats for $t=0,1,2,\dots$ until the task is completed or a predefined termination condition is met. We report the action space of LightAgent in Appendix A.4.
@@ 2.2 Device-Cloud Collaborative Agent System
Figure 4: GUI Performance: On-Device Models vs. Cloud Models.
@@ 2.2.1 Collaborative Control Framework
Task Complexity Assessment. Before task execution begins, LightAgent estimates the difficulty of the task using aggregated historical performance data of the on-device model. An assessment function $f_{\text{assess}}$ analyzes the task description and context to determine two key parameters: the step $\tau$ at which monitoring should begin, and the monitoring interval $\omega$. This preemptive configuration allows the on-device agent to fully utilize its capability without premature—and costly—cloud intervention.
Dynamic Orchestration Policy. During task execution, the system monitors the agent’s behavior once the step counter reaches $\tau$ and at every $\omega$ steps thereafter. A switching function $\mathcal{F}_{\text{switch}}$ evaluates the current GUI state and execution history—including previous actions, state transitions, and task progress—against three criteria: (1) presence of repetitive action patterns, (2) deviation from the expected task trajectory, or (3) inadequate action quality. If any criterion is met, the system switches the current agent from the on-device model $\mathcal{M}_{\text{device}}$ to the cloud model $\mathcal{M}_{\text{cloud}}$, and no further monitoring is performed. This mechanism minimizes unnecessary cloud calls while ensuring reliability through conditional fallback to a more capable cloud model.
The instruction templates for these two components are reported in Appendices A.2.2 and A.2.3.
@@ 2.2.2 Integrated Execution Flow
Input: Task tautau, On-device agent $\mathcal{M}_{\text{device}}$, Cloud agent $\mathcal{M}_{\text{cloud}}$
@@
$(\gamma,\omega)\leftarrow\mathcal{F}_{\text{assess}}(T,\mathcal{M}_{\text{device}})$;
@@ 12 $t\leftarrow t+1$;
2.3 Lightweight MLLM Tuning for On-Device Agents
@@ 2.3.2 Two-Stage Training Protocol
Model training has two stages: supervised fine-tuning (SFT) followed by group relative policy optimization (GRPO) (Shao et al., 2024). In the first stage, chain-of-thought annotations in synthetic training data impart basic reasoning skills and foundational GUI task competence to the small MLLM. This supervised grounding generates meaningful intermediate behaviors, preventing the subsequent reinforcement learning stage from facing overly sparse or uninformative rewards. The second stage is a reinforcement-style policy optimization phase: well-designed reward functions here directly boost the correctness of the model’s output actions and align its behavior with GUI task completion goals.
@@
(ii) Format Reward. $\mathcal{R}_{\text{format}}$ provides a base reward of $\mathcal{R}_{\text{fmt}}\cdot\frac{k}{3}$ based on adherence to Figure 3’s three-block structure, where $k$ measures the degree of conformity (full reward requires complete adherence with $k=3$), and applies a multiplicative penalty $\gamma^{c}$ with coefficient $\gamma<1$ based on the amount of content $c$ outside the template, a mechanism specifically designed to mitigate irrelevant generation common in small-scale MLLM training.
GRPO Training. GRPO eliminates the need for additional value function approximation, as seen in PPO (Schulman et al., 2017), and instead utilizes the average reward from multiple sampled outputs generated in response to the same question as its baseline. Specifically, for each question $q\sim\mathcal{Q}$, a group of outputs $\{o_{1},o_{2},\ldots,o_{\mathcal{G}}\}$ is sampled from the old policy $\pi_{\text{old}}$. The model is optimized by maximizing the following objective:
$\begin{split}J(\theta)=\mathbb{E}_{q\sim\mathcal{Q}}\Bigg[&\mathbb{E}_{{o_{1},\ldots,o_{G}}\sim\pi_{\theta}(\cdot\mid q)}\\ &\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\Big(\rho_{i}(\theta)A_{i},\text{clip}\big(\rho_{i}(\theta),1-\epsilon,1+\epsilon\big)A_{i}\Big)\Bigg]-\beta D_{\text{KL}}\big(\pi_{\theta}\parallel\pi_{\text{ref}}\big)\Bigg].\end{split}$ (5)
@@
$A_{i}=\dfrac{r_{i}-\frac{1}{G}\sum_{j=1}^{G}r_{j}}{\sqrt{\frac{1}{G}\sum_{j=1}^{G}\left(r_{j}-\frac{1}{G}\sum_{k=1}^{G}r_{k}\right)^{2}}}.$ (6)
GRPO’s group-relative approach to calculating advantages aligns seamlessly with the comparative nature of reward models—usually trained on datasets with output comparisons for the same question. Additionally, GRPO incorporates regularization by directly adding the KL divergence (between the trained and reference policies) to the loss function. The KL divergence loss used here follows an unbiased estimator (Hershey & Olsen, 2007):
$D_{\text{KL}}\big(\pi_{\theta}\parallel\pi_{\text{ref}}\big)=\mathbb{E}_{o\sim\pi_{\theta}(\cdot\mid q)}\left[\frac{\pi_{\text{ref}}(o\mid q)}{\pi_{\theta}(o\mid q)}-\log\frac{\pi_{\text{ref}}(o\mid q)}{\pi_{\theta}(o\mid q)}-1\right]$ (7)
@@ 3.1 Experimental Setup
In line with existing work on mobile GUI agents (Liu et al., 2024), we evaluate LightAgent on the academic benchmark AndroidLab (Xu et al., 2024), and further collect four common AndroidLab-based mobile apps for evaluation.
AndroidLab. AndroidLab is an online GUI-task evaluation benchmark built on the Android platform. It comprises nine commonly used applications and 138 evaluation tasks, and supports two input modes: XML mode (Xing et al., 2024) and SoM (Set-of-Mark) mode (Yang et al., 2023). XML mode is tailored to text-only models, where the LLM selects target UI elements directly from the XML representation. SoM mode is intended for multimodal models and uses the Set-of-Mark method: each clickable or focusable element is assigned a unique index, and the LLM specifies elements by their index when issuing operations. Given the growing predominance of MLLMs for GUI tasks, LightAgent primarily adopts the SoM mode; accordingly, all experimental results reported are obtained using SoM mode.
Additional Frequently Used Applications. Existing academic GUI benchmarks, limited by factors such as reproducibility, often neglect to test many of today’s most widely used mobile applications. To offer a more comprehensive evaluation of LightAgent, we augmented AndroidLab with four commonly used mobile apps, contributing a total of 25 tasks. The four apps included are Gmail, Chrome, Reddit, and TikTok. We report a part of designed tasks for these applications in Appendix A.5.

### v2

# OpenPhone: Mobile Agentic Foundation Models

@@ Abstract
With the advancement of multimodal large language models (MLLMs), building GUI agent systems has become an increasingly promising direction—especially for mobile platforms, given their rich app ecosystems and intuitive touch interactions. Yet mobile GUI agents face a critical dilemma: truly on-device models (4B or smaller) lack sufficient performance, while capable models (starting from 7B) are either too large for mobile deployment or prohibitively costly (e.g., cloud-only closed-source MLLMs). To resolve this, we propose OpenPhone, a mobile GUI agent system that leverages device-cloud collaboration to tap the cost-efficiency of on-device models and the high capability of cloud models, while avoiding their drawbacks. Specifically, OpenPhone enhances Qwen2.5-VL-3B via two-stage SFT→GRPO training on synthetic GUI data for strong decision-making, integrates an efficient long-reasoning and memory management mechanism to utilize historical interactions under tight resources, and defaults to on-device execution—only escalating challenging subtasks to the cloud via real-time complexity assessment. Experiments on the online AndroidLab benchmark and diverse apps show OpenPhone matches or nears larger models, with a significant reduction in cloud costs. We have made our OpenPhone available at: https://github.com/HKUDS/OpenPhone.
@@ 1 Introduction
The growing capability of multimodal large language models (MLLMs) enables AI agents to perceive and act within visual environments, particularly through Graphical User Interfaces (GUIs) [31, 22]. Mobile platforms offer a promising domain for this technology for two reasons: first, their vast app ecosystems provide a realistic and diverse testbed, and second, their touchscreen interactions are limited to an intuitive set of primitives, resulting in a compact action space. Despite these advantages, mobile platforms introduce distinct challenges, chiefly severe computational and memory limitations. In light of these factors, our goal is to develop an effective mobile GUI agent, viewing it as a practical milestone toward general-purpose AI.
Current research falls broadly into two groups. The first involves targeted training of open-source MLLMs specifically for GUI-related tasks [23, 8, 19]. These models are relatively compact in size and have achieved notable progress; for instance, UI-Tars-7B [23] outperforms the larger Qwen2.5-VL-32B [5] on mobile GUI tasks. The second approach leverages general-purpose closed-source MLLMs by constructing multi-agent systems and designing well-structured execution pipelines. Thanks to the powerful multimodal comprehension capabilities of state-of-the-art (SOTA) closed-source MLLMs—such as GPT-5 [21], Claude-Sonnet-4 [3], and Gemini-2.5-Pro [7]—their performance on mobile GUI tasks can even surpass that of models specifically trained for such tasks. However, there is no free lunch. Although the performance of models like UI-Tars-7B is impressive given their 7B scale, MLLMs of this size still impose a prohibitive computational burden on contemporary smartphones [15]. A more practical 2B–3B scale, by contrast, typically yields MLLMs with limited capabilities [18]. On the other hand, multi-agent systems based on advanced closed-source MLLMs are plagued by high costs. SOTA closed-source MLLMs are often expensive, and spending several to dozens of dollars to complete a single mobile task is financially impractical.
@@
To answer to the aforementioned questions, we propose a lightweight device-cloud collaborative mobile GUI agent framework—OpenPhone. Specifically, for Question 1, based on a lightweight open-source MLLM (i.e., Qwen2.5-VL-3B), we employ a synthetic data generation pipeline and conduct two-stage training comprising supervised fine-tuning (SFT) and group relative policy optimization (GRPO), resulting in a compact yet powerful on-device GUI agent that achieves performance comparable to larger-scale models. Concurrently, we design an efficient long-reasoning GUI agent paradigm, which is capable of effectively processing historical operation information and endowing the agent with reasoning capabilities to further enhance its performance.
@@
  • Device-cloud collaborative agent system. We propose a device-cloud collaborative agent system that dynamically orchestrates tasks between on-device and cloud models via real-time complexity assessment. This mechanism enables seamless switching to the cloud only when necessary, compensating for on-device limitations, achieving an optimal balance between high performance and reduced operational costs.
  • Comprehensive Evaluation. We evaluate OpenPhone through online experiments on the Android‑based AndroidLab benchmark, complemented by dozens of custom tasks across popular apps to assess real‑world performance. Furthermore, we conduct extensive ablation studies and investigate the impact of different tuning strategies on the lightweight MLLM’s capabilities.
2 The OpenPhone Framework
To effectively address mobile GUI agent tasks, we propose the OpenPhone framework, which has three key modules. Section 2.1 covers strategies to mitigate compact MLLMs’ challenges in GUI tasks—limited capacity and constrained context length. Section 2.2 details a device-cloud collaborative agent system that dynamically schedules on-device and cloud models by task difficulty, balancing cloud resource consumption with GUI task completion rate. Section 2.3 outlines ways to fully leverage limited training data to maximize small MLLMs’ performance gains on GUI tasks.
Figure 2: Overall framework of the proposed OpenPhone.
@@ 2.1 Efficient Reasoning GUI Agent
The on-device GUI agent faces two key challenges. First, mobile devices’ limited computing power necessitates small-sized MLLMs (e.g., Qwen2.5-VL-3B [5]), which currently lack sufficient performance for mobile GUI tasks. To tackle this, OpenPhone enhances the GUI agent with extended chain-of-thought (CoT) reasoning [29] during testing, using test-time scaling laws [26] to boost its capability. Second, limited on-device resources impede long-context handling: high-resolution GUI images take up much available context length, and managing the agent’s execution history is challenging. To alleviate this, OpenPhone uses an efficient text-based summarization scheme—compressing each step’s state into compact textual representations—to support the agent’s long historical context. The output format template is presented in Figure 3, and the detailed instruction template is provided in Appendix A.5.1.
2.1.1 Long-Horizon Reasoning Enhancement
Mobile GUI tasks are often difficult and complex, and humans also use step-by-step reasoning for such operations. Motivated by the success of CoT reasoning and test-time scaling laws, it is natural to apply similar long-form reasoning enhancements to GUI agents. Specifically, the GUI agent’s reasoning—encompassing all factors for task completion—follows a multi-step process: First, it analyzes actionable elements on the current interface and, using historical data, assesses if prior actions met their goals. Next, it evaluates progress toward the user’s task goal and identifies necessary follow-up actions. Lastly, it selects from available functions and parameters to generate the final output. Notably, if historical data shows prior operations failed to yield expected results, the model proactively reflects and adjusts its approach—avoiding repeated errors and potential loops.
2.1.2 Efficient Memory Management
@@
To address this, OpenPhone uses a textual summarization approach: at each step, it distills all information relevant to future actions. As shown in the <STATE_ASSESSMENT> field of Figure 3, these summaries include the current interface state, task progress, the agent’s inferred next action, expected post-action outcome, and potential issues. Much of this content comes from condensing the <REASONING> section—critical for effective stepwise reasoning, especially reflective error correction. Crucially, textual summaries use far fewer tokens than images, enabling long-term history retention (e.g., 10–20 steps) as contextual input, even in resource-constrained mobile environments.
@@ 2.1.3 Overall Process
with $\circ$ denoting sequence concatenation and $h_{0}=\epsilon$. The process repeats for $t=0,1,2,\dots$ until the task is completed or a predefined termination condition is met. We report the action space of OpenPhone in Appendix A.3.
@@ 2.2.1 Collaborative Control Framework
Task Complexity Assessment. Before task execution begins, OpenPhone estimates the difficulty of the task using aggregated historical performance data of the on-device model. An assessment function $f_{\text{assess}}$ analyzes the task description and context to determine two key parameters: the step $\gamma$ at which monitoring should begin, and the monitoring interval $\omega$. This preemptive configuration allows the device agent to fully utilize its capability without premature—and costly—cloud intervention.
Dynamic Orchestration Policy. During task execution, the system monitors the agent’s behavior once the step counter reaches $\gamma$ and at every $\omega$ steps thereafter. A switching function $\mathcal{F}_{\text{switch}}$ evaluates the current GUI state and execution history—including previous actions, state transitions, and task progress—against three criteria: (1) presence of repetitive action patterns, (2) deviation from the expected task trajectory, or (3) inadequate action quality. If any criterion is met, the system switches the current agent from the on-device model $\mathcal{M}_{\text{device}}$ to the cloud model $\mathcal{M}_{\text{cloud}}$, and no further monitoring is performed. This mechanism minimizes unnecessary cloud calls while ensuring reliability through conditional fallback to a more capable cloud model.
The instruction templates for these two components are reported in Appendices A.5.2 and A.5.3.
@@ 2.2.2 Integrated Execution Flow
Input: Task $\tau$, On-device agent $\mathcal{M}_{\text{device}}$, Cloud agent $\mathcal{M}_{\text{cloud}}$
@@
$(\gamma,\omega)\leftarrow\mathcal{F}_{\text{assess}}(\tau,\mathcal{M}_{\text{device}})$;
@@ 12  $t\leftarrow t+1$;
2.3 Lightweight MLLM Tuning
@@ 2.3.2 Two-Stage Training Protocol
Model training has two stages: supervised fine-tuning (SFT) followed by group relative policy optimization (GRPO) [25]. In the first stage, chain-of-thought annotations in synthetic training data impart basic reasoning skills and foundational GUI task competence to the small MLLM. This supervised grounding generates meaningful intermediate behaviors, preventing the subsequent reinforcement learning stage from facing overly sparse or uninformative rewards. The second stage is a reinforcement-style policy optimization phase: well-designed reward functions here directly boost the correctness of the model’s output actions and align its behavior with GUI task completion goals.
@@
$R_{acc}$ and $R_{fmt}$ denote the rewards for answer accuracy and formatting correctness, respectively, during reinforcement learning. By default, both values are set to $1$.
@@
(ii) Format Reward. $\mathcal{R}_{\text{format}}$ provides a base reward of $\mathcal{R}_{\text{fmt}}\cdot\frac{k}{3}$ based on adherence to Figure 3’s three-block structure, where $k$ measures the degree of conformity (full reward requires complete adherence with $k=3$), and applies a multiplicative penalty $\psi^{c}$ with coefficient $\psi<1$ based on the amount of content $c$ outside the template, a mechanism specifically designed to mitigate irrelevant generation common in small-scale MLLM training.
GRPO Training. GRPO eliminates the need for additional value function approximation, as seen in PPO [24], and instead utilizes the average reward from multiple sampled outputs generated in response to the same question as its baseline. Specifically, for each question $q\sim\mathcal{Q}$, a group of outputs $\{o_{1},o_{2},\ldots,o_{\mathcal{G}}\}$ is sampled from the old policy $\pi_{\text{old}}$. The model is optimized by maximizing the following objective:
$\small\begin{split}J(\theta)=\mathbb{E}_{q\sim\mathcal{Q}}\Bigg[\mathbb{E}_{{o_{1},\ldots,o_{G}}\sim\pi_{\theta}(\cdot\mid q)}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\min\Big(\rho_{i}(\theta)A_{i},\text{clip}\big(\rho_{i}(\theta),1-\epsilon,1+\epsilon\big)A_{i}\Big)\Bigg]-\beta D_{\text{KL}}\big(\pi_{\theta}\parallel\pi_{\text{ref}}\big)\Bigg].\end{split}$ (5)
@@
$\small A_{i}=\dfrac{r_{i}-\frac{1}{G}\sum_{j=1}^{G}r_{j}}{\sqrt{\frac{1}{G}\sum_{j=1}^{G}\left(r_{j}-\frac{1}{G}\sum_{k=1}^{G}r_{k}\right)^{2}}}.$ (6)
GRPO’s group-relative approach to calculating advantages aligns seamlessly with the comparative nature of reward models—usually trained on datasets with output comparisons for the same question. Additionally, GRPO incorporates regularization by directly adding the KL divergence (between the trained and reference policies) to the loss function. The KL divergence loss used here follows an unbiased estimator [11]:
$D_{\text{KL}}\big(\pi_{\theta}\parallel\pi_{\text{ref}}\big)=\mathbb{E}_{o\sim\pi_{\theta}(\cdot\mid q)}\Bigl[\frac{\pi_{\text{ref}}(o\mid q)}{\pi_{\theta}(o\mid q)}-\log\frac{\pi_{\text{ref}}(o\mid q)}{\pi_{\theta}(o\mid q)}-1\Bigr]$ (7)
@@ 3.1 Experimental Setup
In line with existing work on mobile GUI agents [19], we evaluate OpenPhone on the academic benchmark AndroidLab [35], and further collect four common AndroidLab-based mobile apps for evaluation. Benchmark details are in Appendix A.2.
Baseline Methods. Comparisons use two primary model groups: (1) General-purpose vision-capable large models: closed-source ones (GPT series [13], Gemini family [7], Claude family), and open-source multimodal models (Qwen-VL family [5, 4], Llama-3.1 [9], GLM series [12]); (2) GUI-specialized models: AutoGLM, AutoGLM-Mobile [34], UI-Tars family [23], V-Droid [8], UI-Genie-Agent [32], MobileUse [17], and two lightly fine-tuned open-source variants (Llama-3.1-8B (ft), GLM-4-9B (ft)).
3.2 Online Agent Evaluation